# Relative Contrastive Learning for Sequential Recommendation with Similarity-based Positive Pair Selection

### v1

@@ 1. Introduction
Sequential Recommendation models (Li et al., 2019; Liu et al., 2021b; Qiu et al., 2019, 2020, 2021b; Zhu et al., 2017; Guo et al., 2018; Rendle et al., 2010; Wang et al., 2015; Wang and Shen, 2022; Qiu et al., 2019, 2020, 2021b; Sun et al., 2019; Kang and McAuley, 2018) predict a user’s subsequent interaction based on their historical sequence, which aims at capturing both short-term preferences and long-term evolving interests. However, sequential recommendation models like GRU4Rec (Hidasi et al., 2015) and SASRec (Kang and McAuley, 2018) encounter challenges of the inherent sparsity and noise within the data. In response, Contrastive Learning (CL) based methods (Qin et al., 2023; Xie et al., 2022; Liu et al., 2021a; Dang et al., 2022), exemplified by recent advancements (Chen et al., 2020; Chen and He, 2021; Hjelm et al., 2018), leverage diverse views to enhance the learned representations of sequences, offering a potential solution to address these data-related limitations in sequential recommendation models.
While CL aims to enhance sequence representations by maximizing agreement among augmented views of the same sequence and distancing views of different sequences, current CL-based methods (Xie et al., 2022; Chong et al., 2023; Hasan et al., 2022b; d’Aquin et al., 2020; Xia et al., 2021; Ding et al., 2022; Qiu et al., 2021a) often rely on instinctual identification of random augmentation operations, like random sequence or model perturbations (‘crop’ and ‘mask’ operations). The process of identifying effective augmentation operations requires specific domain knowledge and meticulous design, potentially causing interference with the original user intent and introducing additional noise during model training (Qiu et al., 2021c; Hasan et al., 2022a).
@@
Recently, supervised contrastive learning (SCL) based methods (Qiu et al., 2021c; Huang et al., 2023; Wang et al., 2023a) suggest choosing same-target sequences (interaction sequences with the same target item) as positive samples. They leverage real sequences instead of augmented sequences and aim at improving the alignment among sequences sharing the same intents. For instance, ContraRec (Wang et al., 2023a) introduces a contrastive learning task to encourage same-target sequences to have similar representations, called context-context contrast. In a similar vein, DuoRec (Qiu et al., 2021c) argues same-target sequences natually contain the same intents and designs a supervised contrastive loss to draw the representations of same-target sequences closer. However, these approaches encounter limitations due to the scarcity of same-target sequences. In the ML-1M and Amazon Beauty datasets, 47.5% and 52.7% of the sequences do not have any same-target sequences, respectively. Consequently, SCL based methods cannot obtain supervised contrast signals on these sequences with no same-target sequences, leading to limited performance improvement.
@@ 3.1. Dual-tiered Positive Sample Selection
In this section, we delve into RCL’s method for selecting positive samples. The SCL-based methods (Huang et al., 2023; Qiu et al., 2021c) have cautioned against the use of data or model-based augmentation in positive sample selection, which could potentially disrupt the inherent user intents. Instead, they try to cluster sequences with identical intents by using the same target sequences as positive samples in contrastive learning paradigm. However, these methods encounter limitations due to the scarcity of same-target sequences. We observe that some negative samples sampled by SCL-based method (Qiu et al., 2021c) in each minibatch actually have high Jaccard or Bigram similarity. Intuitively, the representations of the sequences with high similarity should be close in the latent feature space. Treating the highly similar sequences as negative samples may incorrectly push the representations away and hence degrade the learned representations. Instead, these similar sequences can be used as positive samples to enlarge the original positive sample set.
@@ 3.1.1. Same-target Strong Positive Sample Selection
where $S_{u,a}$ is the sequence sharing the same target item with $S_{u}$. Following the SCL-based methods (Qiu et al., 2021c; Huang et al., 2023), RCL randomly picks one sequence $S_{u,a}$ from $\mathcal{SP}_{u}$ per iteration if $\mathcal{SP}_{u}$ is not empty.
@@ 3.1.2. Similarity-based Weak Positive Pair Selection
    where $S_{u_{1/2}}^{\prime}$ are the string $S_{u_{1/2}}$ with their last characters removed. $m$ and $n$ are the lengths of $S_{u_{1/2}}$. $c(x,y)$ is a function that returns 00 if characters $x$ and $y$ are the same, and $1$ otherwise.
@@ 3.2. Relative Contrastive Learning
where $\mathbf{h}_{u}$,$\mathbf{h}_{a}$, and $\mathbf{h}_{b}$ are the representations of sequences $S_{u}$, $S_{u,a}$, and $S_{u,b}$, respectively.
@@
(12) $\mathcal{L}^{\text{pair}}(S_{u},S_{u,a})=-\log\frac{\exp\left(\mathbf{h}_{u}\cdot\mathbf{h}_{a}/\tau\right)}{\sum_{S_{c}\in A\backslash\{S_{u}\}}\exp\left(\mathbf{h}_{u}\cdot\mathbf{h}_{c}/\tau\right)},$
@@
In SCL methods (Qiu et al., 2021c; Wang et al., 2023a), augmentation is still necessary in case there are no positive samples with the same target item. However, in our framework, this scenario can be prevented.
@@ 3.2.1. Similarity-Based Re-weighting
(16) $\mathcal{L}^{\text{weighted}}(S_{u},S_{u,a})=-\log\frac{s_{u,a}\exp\left(\mathbf{h}_{u}\cdot\mathbf{h}_{a}/\tau\right)}{\sum_{S_{c}\in A\backslash\{S_{u}\}}s_{u,c}\exp\left(\mathbf{h}_{u}\cdot\mathbf{h}_{c}/\tau\right)},$
@@
Input: Training dataset $\left\{S_{u}\right\}_{u=1}^{|\mathcal{U}|}$, hyper-parameters $\lambda$,$\alpha$
@@ 4.1.1. Datasets
  • Yelp [^4] https://www.yelp.com/dataset is widely used for business recommendation. Similar to previous works (Qiu et al., 2021c; Qin et al., 2023), the interaction records after Jan. 1st, 2019 are used in our experiments.
@@
We follow the settings of previous works (Qiu et al., 2021c; Qin et al., 2023; Huang et al., 2023). All interactions are treated as implicit feedback, filtering out users or items appearing fewer than five times. The maximum sequence length for the ML-1M/20M datasets is set to 200, whereas for the other four datasets, the maximum sequence length is set to 50.
For evaluation, we use Top-$K$ Hit Ratio (HR@K) and Top-$K$ Normalized Discounted Cumulative Gain (NDCG@K) across $K$ values of $\{5,10\}$, evaluating rankings across the entire item set for fair comparison, following established methodologies (Qiu et al., 2021c).
@@ 4.1.2. Baselines
  • DuoRec (III) (Qiu et al., 2021c) utilizes both sequences with same target item and model-level augmented sequences as positive samples for contrastive learning.
@@ 4.1.3. Implementation
RCL and all baselines use SASRec as the base model. The embedding size and hidden size are set to 64. The numbers of layers and heads in the Transformer are set to 2. The Dropout (Qiu et al., 2021c) rate on the embedding matrix and the Transformer module is chosen from $\{0.1,0.2,0.3,0.4,0.5\}$. The training batch size is set to 256. We use the Adam (Kang and McAuley, 2018) optimizer with the learning 0.001. $\lambda$ in Equation (18) is chosen from $\{0.1,0.2,0.3,0.4,0.5\}$. $\alpha$ in Section 3.1.2 is chosen from $\{0.025,0.05,0.1,0.15,0.2,0.25\}$. Temperature $\tau$ in Equation (12) is chosen from $\{0.01,0.05,0.1,0.5,1,5\}$.
@@ 4.3. RQ2: Ablation Study
  • Strong solely relies on strong positive samples with $\mathcal{L}^{pair}$ from Eq. (12). In cases where same-target sequences are absent, augmented user representation serves as positive samples instead (Qiu et al., 2021c).
@@
    For sequence $S_{u}$, $\mathcal{L}^{\text{S-W}}_{u}$ treats same-target/similar sequences as positive/negative samples to encourage same-target sequences to have closer representations with $S_{u}$ than similar sequences. $\mathcal{L}^{\text{pair}}(S_{u},S_{u,a})$/$\mathcal{L}^{\text{pair}}(S_{u},S_{u,b})$ treat same-target/similar sequences as positive samples and treat the other sequences except $S_{u}$ in the minibatch as negative samples.
@@ 5.1. Sequential Recommendation
Existing sequential recommendation models (Zhu et al., 2017; Guo et al., 2018; Rendle et al., 2010; Wang et al., 2015; Wang and Shen, 2023; Wang et al., 2023b) mainly rely on recurrent neural networks like GRU (Hidasi et al., 2015) or attention mechanisms such as transformer  (Kang and McAuley, 2018; Sun et al., 2019; d’Aquin et al., 2020; Wang and Shen, 2024) as sequence encoders. Some works  (Qiu et al., 2019, 2020, 2021b) also incorporate graph neural networks to model sequences. Most models  (Guo et al., 2018; Kang and McAuley, 2018) center on the next-item prediction task, which is inherently suitable for predicting subsequent items. The supervision of next-item prediction is limited to data sparsity issue  (Qiu et al., 2021a) and auxiliary tasks such as masked prediction have a semantic gap to the recommendation task.
@@ 5.2. Contrastive Learning
Contrastive learning (CL) has found extensive applications (Oord et al., 2018; Hjelm et al., 2018; He et al., 2020; Chen et al., 2020; Chen and He, 2021) in deep learning to learn latent representations. Based on whether labels are used when generating positive samples, CL can be grouped into self-supervised CL and supervised CL. Some works (d’Aquin et al., 2020; Xie et al., 2022; Xia et al., 2021; Ding et al., 2022; Qiu et al., 2021a) apply self-supervised CL for sequential recommendation. For example, CoSeRec (Liu et al., 2021a) and TiCoSeRec (Dang et al., 2022) perform item-level CL based on model-based augmentations. Recently, some works(Wang et al., 2023a; Qiu et al., 2021c; Huang et al., 2023) propose to apply supervised contrastive learning by taking the sequences with the same target item as positive samples and maximize their similarity. Different from existing methods, we focus on using similar sequences as supplements to positive samples and provide extra training signals for optimizing sequential recommendation models.

### v3

@@ 1. Introduction
Sequential Recommendation models (Li et al., 2019; Liu et al., 2021a; Qiu et al., 2019, 2020, 2021a; Zhu et al., 2017; Guo et al., 2018; Rendle et al., 2010; Wang et al., 2015; Wang and Shen, 2022; Qiu et al., 2019, 2020, 2021a; Sun et al., 2019; Kang and McAuley, 2018) predict a user’s subsequent interaction based on their historical sequence, which aims at capturing both short-term preferences and long-term evolving interests. However, sequential recommendation models like GRU4Rec (Hidasi et al., 2015) and SASRec (Kang and McAuley, 2018) encounter challenges of the inherent sparsity and noise within the data. In response, Contrastive Learning (CL) based methods (Qin et al., 2023; Xie et al., 2022; Liu et al., 2021b; Dang et al., 2022), exemplified by recent advancements (Chen et al., 2020; Chen and He, 2021; Hjelm et al., 2018), leverage diverse views to enhance the learned representations of sequences, offering a potential solution to address these data-related limitations in sequential recommendation models.
While CL aims to enhance sequence representations by maximizing agreement among augmented views of the same sequence and distancing views of different sequences, current CL-based methods (Xie et al., 2022; Chong et al., 2023; Hasan et al., 2022b; d’Aquin et al., 2020; Xia et al., 2021; Ding et al., 2022; Qiu et al., 2021c) often rely on instinctual identification of random augmentation operations, like random sequence or model perturbations (‘crop’ and ‘mask’ operations). The process of identifying effective augmentation operations requires specific domain knowledge and meticulous design, potentially causing interference with the original user intent and introducing additional noise during model training (Qiu et al., 2021b; Hasan et al., 2022a).
@@
Recently, supervised contrastive learning (SCL) based methods (Qiu et al., 2021b; Huang et al., 2023; Wang et al., 2023a) suggest choosing same-target sequences (interaction sequences with the same target item) as positive samples. They leverage real sequences instead of augmented sequences and aim at improving the alignment among sequences sharing the same intents. For instance, ContraRec (Wang et al., 2023a) introduces a contrastive learning task to encourage same-target sequences to have similar representations, called context-context contrast. In a similar vein, DuoRec (Qiu et al., 2021b) argues same-target sequences natually contain the same intents and designs a supervised contrastive loss to draw the representations of same-target sequences closer. However, these approaches encounter limitations due to the scarcity of same-target sequences. In the ML-1M and Amazon Beauty datasets, 47.5% and 52.7% of the sequences do not have any same-target sequences, respectively. Consequently, SCL based methods cannot obtain supervised contrast signals on these sequences with no same-target sequences, leading to limited performance improvement.
@@ 3.1. Dual-tiered Positive Sample Selection
In this section, we delve into RCL’s method for selecting positive samples. The SCL-based methods (Huang et al., 2023; Qiu et al., 2021b) have cautioned against the use of data or model-based augmentation in positive sample selection, which could potentially disrupt the inherent user intents. Instead, they try to cluster sequences with identical intents by using the same target sequences as positive samples in contrastive learning paradigm. However, these methods encounter limitations due to the scarcity of same-target sequences. We observe that some negative samples sampled by SCL-based method (Qiu et al., 2021b) in each minibatch actually have high Jaccard or Bigram similarity. Intuitively, the representations of the sequences with high similarity should be close in the latent feature space. Treating the highly similar sequences as negative samples may incorrectly push the representations away and hence degrade the learned representations. Instead, these similar sequences can be used as positive samples to enlarge the original positive sample set.
@@ 3.1.1. Same-target Strong Positive Sample Selection
where $S_{u,a}$ is the sequence sharing the same target item with $S_{u}$. Following the SCL-based methods (Qiu et al., 2021b; Huang et al., 2023), RCL randomly picks one sequence $S_{u,a}$ from $\mathcal{SP}_{u}$ per iteration if $\mathcal{SP}_{u}$ is not empty.
@@ 3.1.2. Similarity-based Weak Positive Pair Selection
    where $S_{u_{1/2}}^{\prime}$ are the string $S_{u_{1/2}}$ with their last characters removed. $m$ and $n$ are the lengths of $S_{u_{1/2}}$. $c(x,y)$ is a function that returns 0 if characters $x$ and $y$ are the same, and $1$ otherwise.
@@ 3.2. Relative Contrastive Learning
where $\mathbf{h}_{u}$$\mathbf{h}_{a}$, and $\mathbf{h}_{b}$ are the representations of sequences $S_{u}$, $S_{u,a}$, and $S_{u,b}$, respectively.
@@
[…]
@@
In SCL methods (Qiu et al., 2021b; Wang et al., 2023a), augmentation is still necessary in case there are no positive samples with the same target item. However, in our framework, this scenario can be prevented.
@@ 3.2.1. Similarity-Based Re-weighting
[…]
@@
Input: Training dataset $\left\{S_{u}\right\}_{u=1}^{|\mathcal{U}|}$, hyper-parameters $\lambda$$\alpha$
@@ 4.1.1. Datasets
  • Yelp [^4] https://www.yelp.com/dataset is widely used for business recommendation. Similar to previous works (Qiu et al., 2021b; Qin et al., 2023), the interaction records after Jan. 1st, 2019 are used in our experiments.
@@
We follow the settings of previous works (Qiu et al., 2021b; Qin et al., 2023; Huang et al., 2023). All interactions are treated as implicit feedback, filtering out users or items appearing fewer than five times. The maximum sequence length for the ML-1M/20M datasets is set to 200, whereas for the other four datasets, the maximum sequence length is set to 50.
For evaluation, we use Top-$K$ Hit Ratio (HR@K) and Top-$K$ Normalized Discounted Cumulative Gain (NDCG@K) across $K$ values of $\{5,10\}$, evaluating rankings across the entire item set for fair comparison, following established methodologies (Qiu et al., 2021b).
@@ 4.1.2. Baselines
  • DuoRec (III) (Qiu et al., 2021b) utilizes both sequences with same target item and model-level augmented sequences as positive samples for contrastive learning.
@@ 4.1.3. Implementation
RCL and all baselines use SASRec as the base model. The embedding size and hidden size are set to 64. The numbers of layers and heads in the Transformer are set to 2. The Dropout (Qiu et al., 2021b) rate on the embedding matrix and the Transformer module is chosen from $\{0.1,0.2,0.3,0.4,0.5\}$. The training batch size is set to 256. We use the Adam (Kang and McAuley, 2018) optimizer with the learning 0.001. $\lambda$ in Equation (18) is chosen from $\{0.1,0.2,0.3,0.4,0.5\}$. $\alpha$ in Section 3.1.2 is chosen from $\{0.025,0.05,0.1,0.15,0.2,0.25\}$. Temperature $\tau$ in Equation (12) is chosen from $\{0.01,0.05,0.1,0.5,1,5\}$.
@@ 4.3. RQ2: Ablation Study
  • Strong solely relies on strong positive samples with $\mathcal{L}^{pair}$ from Eq. (12). In cases where same-target sequences are absent, augmented user representation serves as positive samples instead (Qiu et al., 2021b).
@@
    For sequence $S_{u}$, $\mathcal{L}^{\text{S-W}}_{u}$ treats same-target/similar sequences as positive/negative samples to encourage same-target sequences to have closer representations with $S_{u}$ than similar sequences. $\mathcal{L}^{\text{pair}}(S_{u},S_{u,a})$$\mathcal{L}^{\text{pair}}(S_{u},S_{u,b})$ treat same-target/similar sequences as positive samples and treat the other sequences except $S_{u}$ in the minibatch as negative samples.
@@ 5.1. Sequential Recommendation
Existing sequential recommendation models (Zhu et al., 2017; Guo et al., 2018; Rendle et al., 2010; Wang et al., 2015; Wang and Shen, 2023; Wang et al., 2023b) mainly rely on recurrent neural networks like GRU (Hidasi et al., 2015) or attention mechanisms such as transformer  (Kang and McAuley, 2018; Sun et al., 2019; d’Aquin et al., 2020; Wang and Shen, 2024) as sequence encoders. Some works  (Qiu et al., 2019, 2020, 2021a) also incorporate graph neural networks to model sequences. Most models  (Guo et al., 2018; Kang and McAuley, 2018) center on the next-item prediction task, which is inherently suitable for predicting subsequent items. The supervision of next-item prediction is limited to data sparsity issue  (Qiu et al., 2021c) and auxiliary tasks such as masked prediction have a semantic gap to the recommendation task.
@@ 5.2. Contrastive Learning
Contrastive learning (CL) has found extensive applications (Oord et al., 2018; Hjelm et al., 2018; He et al., 2020; Chen et al., 2020; Chen and He, 2021) in deep learning to learn latent representations. Based on whether labels are used when generating positive samples, CL can be grouped into self-supervised CL and supervised CL. Some works (d’Aquin et al., 2020; Xie et al., 2022; Xia et al., 2021; Ding et al., 2022; Qiu et al., 2021c) apply self-supervised CL for sequential recommendation. For example, CoSeRec (Liu et al., 2021b) and TiCoSeRec (Dang et al., 2022) perform item-level CL based on model-based augmentations. Recently, some works(Wang et al., 2023a; Qiu et al., 2021b; Huang et al., 2023) propose to apply supervised contrastive learning by taking the sequences with the same target item as positive samples and maximize their similarity. Different from existing methods, we focus on using similar sequences as supplements to positive samples and provide extra training signals for optimizing sequential recommendation models.